HIGH RESOLUTION INFRARED EMISSION SPECTROSCOPY OF DICHLOROBORANE BHCl_{2}

\underline{TSUYOSHI HIRAO}, PINA COLARUSSO and PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1.

The absorption spectrum of dichloroborane BHCl_{2} was accidentally recorded during Fourier transform emission measurements of HBO. Gaseous dichloroborane was generated by the reaction of amorphous boron powder with CaCl_{2} at 1400 C. The bands centred near 2617, 1089 and 892 cm^{−1} were assigned as the ν_{1} (B-H stretch), the ν_{5} (HBCl bend), and the ν_{6} (BCl_{2} antisymmetric stretch) fundamental modes, respectively. Molecular constants were determined from rotational analyses of these bands. This is a first analysis of the rotational structure of the ν_{5} and ν_{6} modes.